# Utilizing Multilingual Encoders to Improve Large Language Models for Low-Resource Languages

### v1

@@ Abstract
Large Language Models (LLMs) excel in English, but their performance degrades significantly on low-resource languages (LRLs) due to English-centric training. While methods like LangBridge align LLMs with multilingual encoders such as the Massively Multilingual Text-to-Text Transfer Transformer (mT5), they typically use only the final encoder layer. We propose a novel architecture that fuses all intermediate layers, enriching the linguistic information passed to the LLM. Our approach features two strategies: (1) a Global Softmax weighting for overall layer importance, and (2) a Transformer Softmax model that learns token-specific weights. The fused representations are mapped into the LLM’s embedding space, enabling it to process multilingual inputs. The model is trained only on English data, without using any parallel or multilingual data. Evaluated on XNLI, IndicXNLI, Sinhala News Classification, and Amazon Reviews, our Transformer Softmax model significantly outperforms the LangBridge baseline. We observe strong performance gains in LRLs, improving Sinhala classification accuracy from 71.66% to 75.86% and achieving clear improvements across Indic languages such as Tamil, Bengali, and Malayalam. These specific gains contribute to an overall boost in average XNLI accuracy from 70.36% to 71.50%. This approach offers a scalable, data-efficient path toward more capable and equitable multilingual LLMs.
@@ I Introduction
To address these limitations, we propose a method that feeds all multilingual encoder layers into the LLM using a learned, weighted fusion by extending LangBridge architecture. Unlike prior work that uses static final-layer outputs [21], our model learns to adaptively combine representations across all layers, capturing diverse linguistic features and improving interpretability.
Our methodology targets multilingual understanding and reasoning tasks where the input is multilingual (English and non-English) and the LLM generated output is in English. We train the model exclusively on English instruction data and evaluate it in a zero-shot [^1] In this context, zero-shot means that the model is trained only on English-language data but is evaluated directly on the same task in languages it has not encountered during training. setting on other languages without any multilingual fine-tuning. This setup reflects real-world scenarios where task-specific instruction data is not available in low-resource languages.
In our experiments, we evaluate the proposed method on a diverse set of datasets covering multiple languages, including XNLI [32] (15 languages), IndicXNLI [33] (12 Indic languages), Sinhala News Classification [34], and Amazon Multilingual Reviews [35] (covering languages such as English, Spanish, French, German, Chinese, and Japanese). These datasets allow us to assess the effectiveness of our approach across both high-resource and low-resource languages including Sinhala, Tamil, Hindi, Malayalam, Assamese, and Odia.
Empirically, we observe that our approach yields consistent improvements across languages, with notably higher gains for languages whose scripts are similar to English. The ability to leverage shared script and structural similarities enables more effective transfer, offering a practical benefit for expanding LLM capabilities to low-resource, script-similar languages without requiring additional annotated data.
@@ II-A Aligning Pretrained Representations
Aligning pretrained encoder outputs with LLMs has gained traction in multi-modal tasks (e.g., vision-language models [27]) and is now being extended to multilingual settings. Approaches such as LangBridge[21] and MindMerger[22] aim to equip English-centric LLMs with multilingual capabilities by incorporating outputs from multilingual encoders.
LangBridge uses a soft replacement strategy, projecting the multilingual encoder’s final-layer hidden states into the LLM’s input space via a learned transformation. MindMerger introduces a two-stage process: (1) a mapping stage using bilingual pairs to align multilingual representations with the LLM, and (2) an augmentation stage to integrate them with the LLM’s inputs.
However, both methods rely solely on the final-layer encoder output, neglecting rich linguistic information present in intermediate layers [26]. Moreover, MindMerger depends on parallel instruction or bilingual data, which is scarce for many low-resource languages, limiting its applicability in zero-shot scenarios.
In contrast, our approach fuses all encoder layers through a learned, weighted combination, requiring only English instruction data. This improves multilingual reasoning without parallel supervision and provides interpretability by revealing which encoder layers contribute most.
@@ II-B Using All Layers of the Encoder
Prior work shows that intermediate encoder layers capture diverse linguistic features useful for downstream tasks. Methods such as weighted summation or attention-based fusion [16] have been applied in English classification and Neural Machine Translation. For example, Bapna et al.[28] use a trainable softmax over encoder layers for neural machine translation, while Liu et al.[29] introduce fixed, learned scaling vectors per layer. DWAtt [30] further improves layer selection through attention mechanism.
These strategies have proven effective in monolingual or bilingual contexts, but have not been extended to multilingual instruction-following LLMs, especially where input is non-English and output is English. Our work bridges this gap by applying deep layer fusion to cross-lingual reasoning in LLMs, without multilingual fine-tuning.
@@ III Methodology
Given a multilingual instruction prompt, we first pass it through a frozen multilingual encoder, which outputs token-level hidden states from all transformer layers. These multi-layered representations encode different linguistic abstractions, from syntactic signals in early layers to semantic and task-relevant features in deeper ones. To effectively leverage this full depth, we introduce a trainable depth-wise weighting mechanism that learns to assign importance scores to each layer per token, generating a single fused representation per token.
These fused token embeddings are then projected into the input space of an English-centric LLM using a learned linear transformation. The encoder and LLM remain frozen during training; however, the fusion and projection components are trained end-to-end using only English instruction–response data. This setup enables the model to align multilingual inputs with English LLM reasoning capabilities without the need for multilingual fine-tuning or parallel instruction data.
@@ III-A Fusion Module
The Fusion Module combines all intermediate hidden states from the multilingual encoder into a single embedding for each token.
For each token, the multilingual encoder produces a sequence of hidden states $h_{1},h_{2},\dots,h_{L}\in\mathbb{R}^{d}$, where $d$ is the embedding dimension. The Fusion Module computes a weighted sum of these representations:
@@ III-A1 Global Softmax Weighting
where base_temp is a fixed constant (e.g., $10^{2}$), factor is a large scaling constant (e.g., $10^{5}$), and temp is a learnable parameter. This formulation ensures a stable yet adaptable temperature that enables sharp or smooth attention over layers depending on task dynamics.
@@ IV-A Datasets
To evaluate the effectiveness of our proposed method, we conduct experiments on a diverse set of multilingual benchmarks that assess natural language understanding, reasoning, and generalization to low-resource languages. We use the XNLI dataset [32], a widely used benchmark for multilingual natural language inference that requires deep understanding and reasoning across 15 languages. To further evaluate performance on low-resource languages, we utilize the IndicXNLI dataset [33], an extension of XNLI designed for Indian languages with limited resources. Additionally, we include the Sinhala News Classification dataset [34], which involves categorizing news headlines written in Sinhala. We also employ the multilingual Amazon Reviews sentiment analysis dataset [35] to test sentiment classification across multiple languages.
@@ IV-B Models and Baseline
Our primary baseline is LangBridge [21], a recent and competitive architecture designed to enhance the multilingual capabilities of LLMs by aligning a multilingual encoder with an English-centric decoder. This was enhanced by Mindmerger  [22] with the use of parallel instruction data. In our experiments, we do not consider using parallel instruction data (non-English instruction and English response) because it does not represent real world scenarios where parallel instruction data are limited in non-English languages. We compare our proposed method against this baseline using two proposed variants of our model: Global Softmax Weighting (Global Softmax) and Token-Wise Softmax Weighting (Transformer SM). These comparisons allow us to assess the impact of fine-grained, token-specific attention over encoder layers in improving multilingual reasoning and understanding.
@@ V-B Improvements in Low-Resource Languages
The most significant gains are observed in LRLs, which are traditionally underserved by English-centric LLMs. In the IndicXNLI benchmark (Table II), both fusion strategies outperform LangBridge in nearly all languages, with marked improvements in Assamese, Gujarati, and Malayalam. Similarly, on the Sinhala News Classification task (Table III), Transformer SM notably boosts performance over LangBridge, indicating its ability to capture meaningful linguistic representations even in limited-resource settings.
These results validate our core hypothesis that incorporating representations from all encoder layers provides richer linguistic grounding, particularly crucial for LRLs where single-layer representations may not be sufficient to convey complex linguistic structure.
@@ V-E Layer Weight Distribution Analysis
To better understand the model’s behavior, we visualize the learned layer weights in Figure 2. The plot reveals that lower layers contribute minimally, while middle and upper layers receive significantly higher weights, peaking around layers 15–19. This pattern is consistent with prior findings that intermediate encoder layers often encode critical syntactic and semantic features [37]. The weight distribution also highlights that the model learns to adaptively attend to deeper layers, supporting the motivation behind our layer fusion strategy.
@@ VI Conclusion
This work introduces a fusion-based method to enhance the multilingual capabilities of English-centric Large Language Models (LLMs) by incorporating all hidden layers of a frozen multilingual encoder. Unlike prior approaches that rely only on the final layer, our method uses either global or token-wise fusion to generate richer token embeddings, trained using a prefix language modeling objective without requiring parallel multilingual data.
Experiments on XNLI, IndicXNLI, Sinhala News Classification, and Amazon Reviews show consistent improvements, especially in low-resource languages. The Transformer Softmax variant achieved the highest gains—for example, improving Sinhala News Classification accuracy from 71.66% to 75.86%. These results are statistically significant and confirm the value of deep encoder fusion in zero-shot multilingual settings.
However, due to computational constraints, the evaluations were limited to models with 1B parameters. It remains to be tested whether the proposed method provides similar gains when applied to larger or already multilingual-tuned models such as Mistral-MoE. This is a key direction for future research.
Our findings highlight the value of multilayer encoder fusion in bridging performance gaps for low-resource languages and offer a scalable path to more inclusive multilingual NLP.

### v2

@@ Abstract
Large Language Models (LLMs) excel in English, but their performance degrades significantly on low-resource languages (LRLs) due to English-centric training. While there are methods to align LLMs with multilingual encoders such as the Massively Multilingual Text-to-Text Transfer Transformer (mT5), they typically use only the final encoder layer. We propose a novel architecture that fuses all intermediate layers, enriching the linguistic information passed to the LLM. Our approach features two strategies: (1) a Global Softmax weighting for overall layer importance, and (2) a Transformer Softmax model that learns token-specific weights. The fused representations are mapped into the LLM’s embedding space, enabling it to process multilingual inputs. The model is trained only on English data, without using any parallel or multilingual data. Evaluated on XNLI, IndicXNLI, Sinhala News Classification, and Amazon Reviews, our Transformer Softmax model significantly outperforms the existing baseline. We observe strong performance gains in LRLs, improving Sinhala classification accuracy from 71.66% to 75.86% and achieving clear improvements across Indic languages such as Tamil, Bengali, and Malayalam. These specific gains contribute to an overall boost in average XNLI accuracy from 70.36% to 71.50%. This approach offers a scalable, data-efficient path toward more capable and equitable multilingual LLMs [^1] Our code is available at https://github.com/ImalshaD/MultiLingualImprove.
@@ I Introduction
To address these limitations, we propose a method that feeds all multilingual encoder layers into the LLM using a learned, weighted fusion by extending LangBridge architecture. Unlike prior work that uses static final-layer outputs [21], our model learns to adaptively combine representations across all layers, capturing diverse linguistic features and improving interpretability. Our methodology targets multilingual understanding and reasoning tasks where the input is multilingual (English and non-English) and the LLM generated output is in English. We train the model exclusively on English instruction data and evaluate it in a zero-shot [^2] In this context, zero-shot means that the model is trained only on English-language data but is evaluated directly on the same task in languages it has not encountered during training. setting on other languages without any multilingual fine-tuning. This setup reflects real-world scenarios where task-specific instruction data is not available in low-resource languages.
In our experiments, we evaluate the proposed method on a diverse set of datasets covering multiple languages, including XNLI [27] (15 languages), IndicXNLI [28] (12 Indic languages), Sinhala News Classification [29, 30], and Amazon Multilingual Reviews [31] (covering languages such as English, Spanish, French, German, Chinese, and Japanese). Empirically, we observe that our approach yields consistent improvements across languages, with notably higher gains for languages whose scripts are similar to English. The ability to leverage shared script and structural similarities enables more effective transfer, offering a practical benefit for expanding LLM capabilities to low-resource, script-similar languages without requiring additional annotated data.
@@ II-A Aligning Pretrained Representations
Aligning pretrained encoder outputs with LLMs has gained traction in multi-modal tasks (e.g., vision-language models [32]) and is now being extended to multilingual settings. Approaches such as LangBridge [21] and MindMerger [22] aim to equip English-centric LLMs with multilingual capabilities by incorporating outputs from multilingual encoders. LangBridge uses a soft replacement strategy, projecting the multilingual encoder’s final-layer hidden states into the LLM’s input space via a learned transformation. MindMerger introduces a two-stage process: (1) a mapping stage using bilingual pairs to align multilingual representations with the LLM, and (2) an augmentation stage to integrate them with the LLM’s inputs. However, both methods rely solely on the final-layer encoder output, neglecting rich linguistic information present in intermediate layers [26]. Moreover, MindMerger depends on parallel instruction or bilingual data, which is scarce for many low-resource languages, limiting its applicability in zero-shot scenarios.
@@ II-B Using All Layers of the Encoder
Prior work shows that intermediate encoder layers capture diverse linguistic features useful for downstream tasks. Methods such as weighted summation or attention-based fusion [16] have been applied in English classification and Neural Machine Translation. For example, Bapna et al.[33] use a trainable softmax over encoder layers for neural machine translation, while Liu et al.[34] introduce fixed, learned scaling vectors per layer. DWAtt [35] further improves layer selection through an attention mechanism. These strategies have proven effective in monolingual or bilingual contexts, but have not been extended to multilingual instruction-following LLMs, especially where input is non-English and output is English.
@@ III Methodology
Given a multilingual instruction prompt, we first pass it through a frozen multilingual encoder, which outputs token-level hidden states from all transformer layers. These multi-layered representations encode different linguistic abstractions, from syntactic signals in early layers to semantic and task-relevant features in deeper ones. To effectively leverage this full depth, we introduce a trainable depth-wise weighting mechanism that learns to assign importance scores to each layer per token, generating a single fused representation per token. These fused token embeddings are then projected into the input space of an English-centric LLM using a learned linear transformation. The encoder and LLM remain frozen during training; however, the fusion and projection components are trained end-to-end using only English instruction–response data. This setup enables the model to align multilingual inputs with English LLM reasoning capabilities without the need for multilingual fine-tuning or parallel instruction data.
@@ III-A Fusion Module
The Fusion Module combines all intermediate hidden states from the multilingual encoder into a single embedding for each token. For each token, the multilingual encoder produces a sequence of hidden states $h_{1},h_{2},\dots,h_{L}\in\mathbb{R}^{d}$, where $d$ is the embedding dimension. The Fusion Module computes a weighted sum of these representations:
@@ III-A1 Global Softmax Weighting
where base_temp is a fixed constant (e.g., $10^{2}$), factor is a large scaling constant (e.g., $10^{5}$), and temp is a learnable parameter.
@@ IV-A Datasets
To evaluate the effectiveness of our proposed method, we conduct experiments on a diverse set of multilingual benchmarks that assess natural language understanding, reasoning, and generalization to low-resource languages. We use the XNLI dataset [27]. To further evaluate performance on low-resource languages, we utilize the IndicXNLI dataset [28], an extension of XNLI designed for Indian languages with limited resources. Additionally, we include the Sinhala News Classification dataset [30], which involves categorizing news headlines written in Sinhala. We also employ the multilingual Amazon Reviews sentiment analysis dataset [31] to test sentiment classification across multiple languages.
@@ IV-B Models and Baseline
Our primary baseline is LangBridge [21], a recent and competitive architecture designed to enhance the multilingual capabilities of LLMs by aligning a multilingual encoder with an English-centric decoder. This was enhanced by Mindmerger [22] with the use of parallel instruction data. In our experiments, we do not consider using parallel instruction data (non-English instruction and English response) because it does not represent real world scenarios where parallel instruction data are limited in non-English languages. Thus Mindmerger is not considered a baseline. We compare our proposed method against LangBridge using two variants of our model: Global Softmax Weighting (Global Softmax) and Token-Wise Softmax Weighting (Transformer SM).
@@ V-B Improvements in Low-Resource Languages
The most significant gains are observed in LRLs, which are traditionally underserved by English-centric LLMs. In the IndicXNLI benchmark (Table II), both fusion strategies outperform LangBridge in nearly all languages, with marked improvements in Assamese, Gujarati, and Malayalam. Similarly, on the Sinhala News Classification task (Table III), Transformer SM notably boosts performance over LangBridge, indicating its ability to capture meaningful linguistic representations even in limited-resource settings. These results validate our core hypothesis that incorporating representations from all encoder layers provides richer linguistic grounding, particularly crucial for LRLs where single-layer representations may not be sufficient to convey complex linguistic structure.
@@ V-E Layer Weight Distribution Analysis
To better understand the model’s behavior, we visualize the learned layer weights in Figure 2. The plot reveals that lower layers contribute minimally, while middle and upper layers receive significantly higher weights, peaking around layers 15–19. This pattern is consistent with prior findings that intermediate encoder layers often encode critical syntactic and semantic features [38].
@@ VI Conclusion
This work introduces a fusion-based method to enhance the multilingual capabilities of English-centric Large Language Models (LLMs) by incorporating all hidden layers of a frozen multilingual encoder. Unlike prior approaches that rely only on the final layer, our method uses either global or token-wise fusion to generate richer token embeddings, trained using a prefix language modeling objective without requiring parallel multilingual data. Experiments on different datasets show consistent improvements, especially in LRLs. The Transformer Softmax variant achieved the highest gains. These results confirm the value of deep encoder fusion in zero-shot multilingual settings. Thus, our findings highlight the value of multilayer encoder fusion in bridging performance gaps for LRLs and offer a scalable path to more inclusive multilingual NLP. However, due to computational constraints, the evaluations were limited to an LLM with 1B parameters. It remains to be tested whether our method provides similar gains when applied to larger or already multilingual-tuned LLMs such as Mistral-MoE. This is a key direction for future research.